### v1

@@ Abstract
Context. Accretion bursts from low-mass young stellar objects (YSOs) are known for many decades. In recent years, the first accretion bursts of massive YSOs (MYSOs) were observed. These phases of intense protostellar growth are of particular importance for studying massive star formation. Bursts of MYSOs are accompanied by flares of Class II methanol masers (hereafter masers), caused by an increase in exciting mid-infrared (MIR) emission. They can lead to long-lasting thermal afterglows of the dust continuum radiation visible at infrared (IR) and (sub)mm wavelengths. Furthermore, they might cause a scattered light echo. The G323.460.08 (hereafter G323) event, which shows all these features, extends the small sample of known MYSO bursts.
@@
Methods. Burst light curves, as well as the pre-burst spectral energy distribution (SED) were established from archival IR data. The properties of the MYSO, including its circumstellar disk and envelope, were derived by using static radiative transfer modeling of pre-burst data. For the first time, TDRT was used to predict the temporal evolution of the SED. Observations with SOFIA/HAWC+ were performed to constrain the burst energy from the strength of the thermal afterglow. Image subtraction and ratioing was applied to reveal the light echo.
Results. The G323 accretion burst is confirmed. It reached its peak in late 2013/early 2014 with a -band increase of 2.5 mag. Both -band and integrated maser flux densities follow an exponential decay. TDRT indicates that the duration of the thermal afterglow in the far-infrared (FIR) can exceed the burst duration by years. The latter was proved by SOFIA observations, which indicate a flux increase of % at and % at m in 2022 (2 years after the burst end). A one-sided light echo emerged that was propagating into the interstellar medium.
Conclusions. The burst origin of the G323 maser flare is verified. TDRT simulations revealed the strong influence of the burst energetics and the local dust distribution on the strength and duration of the afterglow. The G323 burst is probably the most energetic MYSO burst observed so far. Within , an energy of was released. The short timescale points to the accretion of a compact body, while the burst energy corresponds to an accumulated mass of at least and possibly even more if the protostar is bloated. In this case, the accretion event might have triggered protostellar pulsations, which give rise to the observed maser periodicity. The associated IR light echo is the second observed from a MYSO burst.
@@ 1 Introduction
Episodic accretion events are phases of strongly enhanced mass accumulation during (proto-)stellar growth. These are not restricted to young stars of low and intermediate mass (Hartmann & Kenyon, 1996; Audard et al., 2014), but also occur during the formation of high-mass stars (Caratti o Garatti et al., 2017; Hunter et al., 2017; Stecklum et al., 2021). The energy released during such accretion bursts strongly affects the MYSO and its circumstellar environment in various ways. Heating of the dust in both the circumstellar disk and envelope leads to a temporary increase in the dust continuum emission. Depending on the strength and duration of the burst, this partially or completely affects the SED. The timescales on which the spectral energy distribution (SED) changes are wavelength dependent, where different wavelength regions trace different spatial regions (e.g., Contreras Peña et al., 2020). Therefore, accretion bursts provide a unique opportunity for ‘screening’ MYSOs, which are often deeply embedded throughout their formation time.
@@
A specific aspect of MYSO accretion bursts is that their high luminosity leads to the sublimation of volatile substances that were trapped in the ice mantles of dust grains. Thus, for certain molecules, such as methanol in particular, maser emission can occur once the excitation conditions, for example, specific column density and strong mid-infrared (MIR) radiation field, are satisfied. For this reason, radiatively pumped masers (Class II, Menten, 1991a; Sobolev et al., 1997a; Cragg et al., 2005) are a very good tracer of MYSOs (Breen et al., 2013). These masers will flare during accretion bursts, which makes them a reliable burst alert. Particularly useful is the 6.7 GHz transition (Menten, 1991b), as this is usually the brightest. During the burst, the maser spots can be relocated, thus providing information on the local structures such as spiral arms in a disk (see, e.g., Burns et al., 2020).
@@
Even after the burst is over, the FIR fluxes will remain elevated for quite some time, as witnessed for the MYSO G358.93-0.03 burst (Stecklum et al., 2021). This thermal afterglow depends on the local dust distribution and is a record of the history of the burst. Its detection represents an a posteriori confirmation for MYSO bursts, that were not or could not be detected by other means. It allows constraining the energy release during the burst, which is fundamental for understanding the triggering mechanisms behind it. Until its recent shutdown, the Stratospheric Observatory for Infrared Astronomy (SOFIA, Erickson & Davidson, 1993; Young et al., 2012) was the only facility offering the capability to verify the increase in FIR flux caused by the burst, thus allowing for the detection of such an afterglow.
@@
The object G323, which is also known as IRAS15254-5621, is a massive star-forming region. It is located at RA: , , J2000. It is covered by various surveys, since it resides in the vicinity of the Galactic center. The red MSX survey (RMS; Lumsden et al., 2013) and the APEX Telescope Large Area Survey of the Galaxy (ATLASGAL; Schuller et al., 2009) revealed the main properties of the region. A bolometric luminosity of was derived by integrating its SED (Lumsden et al., 2013). G323 is accompanied by the compact ATLASGAL clump AGAL 323.459-00.079 (Urquhart et al., 2014), considered a massive cluster progenitor (Csengeri et al., 2017) with a mass of that hosts the MYSO. For this reason, the source is included in the ALMAGAL survey, a large program on ALMA (2019.1.00195.L, PI: S. Molinari), dedicated to studying the evolution of high-mass protocluster formation in the Galaxy. Araya et al. (2005) found blue- and redshifted CS (3–2) as well as 13CO (2–1) line wings, indicative of a molecular outflow. A similar signature in the 18CO (2–1) line is observed by the SEDIGISM survey (Yang et al., 2022). However, Guerra-Varas et al. (2023) did not detect an outflow in the SiO (2-1) line and wings of the HCO+ (2-1) line with the APEX telescope. Recently, Ma et al. (2023) identified this clump as a hub for star formation fed by three filaments. For the hub junction, they derived an extent of 2.4 pc 2.4 pc and a mass of .
@@
Associated weak radio continuum emission was first reported by Haynes et al. (1978) and later mapped by Walsh et al. (1998) who did not resolve the source (). Araya et al. (2005) detected broad radio recombination lines towards this region, indicating the presence of a hypercompact Hii region (HCHii). Similar measurements of Murphy et al. (2010), Kim et al. (2018) confirmed this finding. Murphy et al. (2010) also estimated the extinction towards G323, based on the depth of the silicate absorption feature in a low-resolution IRAS spectrum, which amounts to mag. The object was observed in the ATOMS survey (ALMA Three-millimeter Observations of Massive Star-forming regions, Liu et al., 2020b) in Autumn 2019. The radio recombination line and the continuum data at 3 mm are used by Zhang et al. (2023) to characterize the HCHii.
@@
The massive star-forming region G323 is associated with methanol, water, and hydroxyl masers. The Class II methanol 6.7 GHz transition was first observed by MacLeod et al. (1992). Monitoring of this transition at Hartebeesthoek Radio Astronomy Observatory (HartRAO) revealed an increase in the total flux density from 20 Jy in 2011 (Green et al., 2015) to 7 500 Jy in 2015 (Proven-Adzri et al., 2019), i.e., by a factor of 430. Interestingly, the decay of the methanol maser, which has been ongoing since then, is characterized by a periodic flux variation with a cycle time of 93 d (MacLeod et al., 2021). The first flare evidence was obtained from the brightening of the 6.035 GHz exOH maser (MacLeod et al., 2021). The discovery of the maser flare raises the question whether it is due to an accretion burst, similar to S255IR-NIRS3 (Caratti o Garatti et al., 2017), NGC6334I-MM1 (Hunter et al., 2017, 2018), and G358 (Brogan et al., 2019; MacLeod et al., 2019; Burns et al., 2020; Stecklum et al., 2021). The source is classified as ‘irregular’ in the recent NEOWISE-based variability study of 6.7 GHz maser sources by Song et al. (2023). Their exclusion of the pre-burst epochs prevents the burst detection.
@@ 2.1 VVV(X) survey imaging
The VISTA Variables in Via Lactea Survey (VVV, Minniti et al., 2010) and its extension VVVX (from 2016 to 2019) are ESO public surveys that were conducted with the 4-m VISTA telescope in the red optical/NIR domain. They target the Galactic bulge and part of the adjacent plane. Images of the G323 region were retrieved from the VISTA Science Archive. VVV(X) obtained -band imaging in all of their observing epochs of the region. Data for the and filters are available only for 2010 and one epoch in 2015. Although this does not suffice to produce light curves, a possible color change resulting from the burst can be studied nevertheless. NIR images show the bright counterpart of G323 embedded in a scattering nebulosity. Photometry was established on VVV(X) images, taking into account the core saturation of its point spread function (PSF) in the -band. The affected pixels were given zero weight when fitting the image profile. This was performed using the MPFit2DPeak function of the IDL Astronomy Users Library (Landsman, 1995), using a tilted Moffat function that is appropriate for PSFs in ground-based images (Moffat, 1969). Images with seeing worse than 35 were discarded. The photometry is given in Table 10 of the Appendix.
@@ 2.3 ISAAC spectroscopy
Long-slit spectroscopy with the Infrared Spectrometer and Array Camera ISAAC was performed on June 11, 2013, at the ESO-VLT UT3 telescope, Paranal, Chile. A slit of 03 width was used, producing a spectral resolution of 8 900, at a position angle (PA) of 438, centered on the source (see Fig. 10). The total integration time was 4 min. Nodding along the slit was performed with a throw of 35″ to allow for sky subtraction. The grating was set at a central wavelength of 2.15 m, offering a spectral bandwidth of 0.122 m, covering a wavelength range from 2.086 to 2.210 m. Data reduction was performed using standard IRAF222IRAF (Image Reduction and Analysis Facility) is distributed by the National Optical Astronomy Observatories, which are operated by AURA, Inc., in cooperative agreement with the National Science Foundation. tasks. Each observation was flat-fielded, sky subtracted, and corrected for the distortion caused by long-slit spectroscopy. The atmospheric response was corrected by dividing each spectrum by a telluric standard star (acquired with the same science settings), normalized to the blackbody function at the stellar temperature, and corrected for any absorption feature intrinsic to the star. Originally, we used the telluric standard star for flux calibration. However, probably due to slit flux loss, the resulting calibrated spectrum is = 6.73, mag, that is, 0.6, mag lower than the VVV photometry around that epoch (see Sect. 3.2). Therefore, for calibration, we adopted the closest -band photometric point (16 June 2013). Wavelength calibration was performed using the many OH lines located in this spectral range, resulting in an average uncertainty of 0.2 Å in the extracted spectra. Three spectra, centered on the source and a few arcseconds NE and SW off the source, were extracted from the spectral image, where both continuum and line emission are present. Continuum emission from the final spectral image was removed with the IRAF task “continuum”. Finally, the radial velocities of the observed lines were calculated using single or multiple Gaussian fits (for multiple components), and corrected for the target velocity with respect to the LSR.
@@ 2.9 ALMA observations
The ALMA Cycle 7 project ALMAGAL aims to observe the 1.3 mm continuum and lines towards dense molecular clumps in the Galactic Plane at a sensitivity level of . Measurements are made with the 12-m array and the 7-m Alma Compact Array (ACA). Data for G323 were retrieved from the ALMA Science Archive (e.g., Stoehr et al., 2014). For this paper, we used the available 12-m array measurements on ALMAGAL field 767784 (observing date: 30-December-2019), where we worked on the data products obtained by the standard pipeline in CASA (Common Astronomy Software Applications package, McMullin et al., 2007). This data set has a typical spatial resolution of 12. In the course of this work, a discrepancy was observed between the flux density at 1.4 mm obtained from the ALMAGAL data and the 3 mm measurement by Zhang et al. (2023). Therefore, the corresponding data set from Zhang et al. (2023), obtained at about the same epoch, was also analyzed using CASA.
@@ 3.3 The scattered light echo
LEs are also detected in the NEOWISE and images, which were retrieved using the ICORE tool (Masci, 2013). This is not self-evident, as the scattering cross-section of interstellar dust grains at longer wavelengths is smaller (e.g., Draine, 2003). However, the lower scattering efficiency is more than compensated by higher flux densities in the NEOWISE bands compared to (cf. Fig. 15). The features of the LEs and their evolution seen by NEOWISE correspond to what is found in the -band.
Near the source, the spatial brightness distribution of the LE echo represents a record of the burst history. Thus, the maximum of the ratio values provides an estimate of the maximum burst strength. This holds for the - and -bands, where the maximum values at various locations are around 20. Further away, the echo strength decreases as a result of spatial propagation and superposition along the line of sight. This is already the case for the -band. Clearly, the LE is not circular-symmetric but extends southeast and is missing in the opposite direction. This indicates a non-uniform dust distribution in the environment of G323. The absence of an LE in the northwest could be related to the curved rim-shaped structure in this area (cf. Fig. 1) that could block or shadow it.
@@
The remote LE is still faintly visible in NEOWISE images taken in 2023 (cf. Fig. 24). Its light curves (Fig. 7) were established by integrating the fluxes in the , , and bands over this region, marked in Fig. 6. The subtracted background was derived from an area north-west of the MYSO, devoid of stars. Notably, these light curves cover the burst rise which is missing in the on-source data (cf. Fig. 4) because of the time gap between the and NEOWISE missions. The primary characteristics that can be drawn from them are the following. The LE peak occurs slightly more than four years after the burst peak. Its amplitude is shallower than that of the burst, which is due to echo propagation. The color change confirms the ‘bluening’ during the burst. Due to the wavelength dependence of the scattering cross section, the pre-burst colors of the LE are bluer than those of the source, as indicated by vs. 0.12 and vs. .
@@
For common cases of LEs, as for example those associated with novae, supernovae, and variable stars (Sugerman, 2007; Jiang et al., 2016; Kervella et al., 2008), a superluminal motion of the echo is present. It is a pure geometrical effect that follows from first principles in the case of single scattering (Nemiroff et al., 2016). Whether a prompt YSO LE is superluminal is questionable, since its environment is quite dusty. Thus, multiple scattering may overwhelm the geometric effect on which the superluminal motion rests. For G323, the apparent propagation velocity can be derived from the projected separation between the echo boundary and the MYSO, as well as the time difference between the date of the burst peak and the epoch of observation. For the -band, the respective quantities are 22″ and one year. Together with a distance of 4.08 kpc, this yields 1.4 . Although this confirms the superluminal motion in the -band, it implies that the apparent velocity is smaller for and . Given the small extent of the echo, its apparent velocity is probably close to sub-luminal.
@@
In the optically thick regions, the photons follow random walk paths, where the effective travel distance is given by the diffusion path length. For random walk processes, the diffusion path length is (assuming isotropic scattering), where is the mean free path and is the number of scattering events. Within a fixed time the photons can travel a path with a ‘real’ length of , with c being the speed of light. The ‘real’ path is the same for all wavelengths (it depends only on the travel time and c), but the interaction free path length (and hence the number of scattering events) depends on the optical depth (and hence the wavelength). When the equations are combined, the diffusion length at a given time and wavelength is . This implies a difference in the diffusion path length (and hence the apparent LE extent) for different wavelengths. With (see e.g., Krieger & Wolf, 2021), the expected extent of the LE in compared to that seen in Z can be written as the inverse ratio of the optical depths . For MRN dust (Mathis et al., 1977), the optical depth ratio between the Z and filters is . This would imply that LE propagation is almost a factor 4 slower in Z than in . However, the density within the cloud core is highly variable, and a measurable slowdown is expected only in the densest regions (close to the protostar) and therefore the effect will be smaller.

### v2

@@ Abstract
Context. Accretion bursts from low-mass young stellar objects (YSOs) have been known for many decades. In recent years, the first accretion bursts of massive YSOs (MYSOs) have been observed. These phases of intense protostellar growth are of particular importance for studying massive star formation. Bursts of MYSOs are accompanied by flares of Class II methanol masers (hereafter masers), which are caused by an increase in exciting mid-infrared (MIR) emission. They can lead to long-lasting thermal afterglows of the dust continuum radiation visible at infrared (IR) and (sub)millimeter (hereafter (sub)mm) wavelengths. Furthermore, they might cause a scattered light echo. The G323.460.08 (hereafter G323) event, which shows all these features, extends the small sample of known MYSO bursts.
@@
Methods. Burst light curves, as well as the pre-burst spectral energy distribution (SED) were established from archival IR data. The properties of the MYSO, including its circumstellar disk and envelope, were derived by using static radiative transfer modeling of pre-burst data. For the first time, TDRT was used to predict the temporal evolution of the SED. Observations with SOFIA/HAWC+ were performed to constrain the burst energy from the strength of the thermal afterglow. Image subtraction and ratioing were applied to reveal the light echo.
Results. The G323 accretion burst is confirmed. It reached its peak in late 2013/early 2014 with a -band increase of 2.5 mag. Both -band and integrated maser flux densities follow an exponential decay. TDRT indicates that the duration of the thermal afterglow in the far-infrared (FIR) can exceed the burst duration by years. The latter was proved by SOFIA observations, which indicate a flux increase of % at and % at m in 2022 (2 years after the burst ended). A one-sided light echo emerged that was propagating into the interstellar medium.
Conclusions. The burst origin of the G323 maser flare has been verified. TDRT simulations revealed the strong influence of the burst energetics and the local dust distribution on the strength and duration of the afterglow. The G323 burst is probably the most energetic MYSO burst that has been observed so far. Within , an energy of was released. The short timescale points to the accretion of a compact body, while the burst energy corresponds to an accumulated mass of at least and possibly even more if the protostar is bloated. In this case, the accretion event might have triggered protostellar pulsations, which give rise to the observed maser periodicity. The associated IR light echo is the second observed from a MYSO burst.
@@ 1 Introduction
Episodic accretion events are phases of strongly enhanced mass accumulation during (proto-)stellar growth. These are not restricted to young stars of low and intermediate mass (Hartmann & Kenyon, 1996; Audard et al., 2014), but they also occur during the formation of high-mass stars (Caratti o Garatti et al., 2017; Hunter et al., 2017; Stecklum et al., 2021). The energy released during such accretion bursts strongly affects the massive young stellar object (MYSO) and its circumstellar environment in various ways. Heating of the dust in both the circumstellar disk and envelope leads to a temporary increase in the dust continuum emission. Depending on the strength and duration of the burst, this partially or completely affects the spectral energy distribution (SED). The timescales on which the SED changes are wavelength dependent, where different wavelength regions trace different spatial regions (e.g., Contreras Peña et al., 2020). Therefore, accretion bursts provide a unique opportunity for ”screening” MYSOs, which are often deeply embedded throughout their formation time.
@@
A specific aspect of MYSO accretion bursts is that their high luminosity leads to the sublimation of volatile substances that were trapped in the ice mantles of dust grains. Thus, for certain molecules, such as methanol in particular, maser emission can occur once the excitation conditions, for example, the specific column density and strong mid-infrared (MIR) radiation field, are satisfied. For this reason, radiatively pumped masers (Class II, Menten, 1991a; Sobolev et al., 1997a; Cragg et al., 2005) are a very good tracer of MYSOs (Breen et al., 2013). These masers will flare during accretion bursts, which makes them a reliable burst alert. Particularly useful is the 6.7 GHz transition (Menten, 1991b), as this is usually the brightest. During the burst, the maser spots can be relocated, thus providing information on the local structures such as spiral arms in a disk (see, e.g., Burns et al., 2020).
@@
Even after the burst is over, the far-infrared (FIR) fluxes remain elevated for quite some time, as witnessed for the MYSO G358.93-0.03 burst (Stecklum et al., 2021). This thermal afterglow depends on the local dust distribution and is a record of the history of the burst. Its detection represents an a posteriori confirmation for MYSO bursts, which were not or could not be detected by other means. It allows for the energy release to be constrained, which is fundamental for understanding the triggering mechanisms behind the burst. Until its recent shutdown, the Stratospheric Observatory for Infrared Astronomy (SOFIA, Erickson & Davidson, 1993; Young et al., 2012) was the only facility offering the capability to verify the increase in FIR flux caused by the burst, thus allowing for the detection of such an afterglow.
@@
The object G323, which is also known as IRAS15254-5621, is a massive star-forming region. It is located at RA: , , J2000. It is covered by various surveys, since it resides in the vicinity of the Galactic center. The red MSX survey (RMS; Lumsden et al., 2013) and the APEX Telescope Large Area Survey of the Galaxy (ATLASGAL; Schuller et al., 2009) revealed the main properties of the region. A bolometric luminosity of was derived by integrating its SED (Lumsden et al., 2013). G323 is accompanied by the compact ATLASGAL clump AGAL 323.459-00.079 (Urquhart et al., 2014), considered a massive cluster progenitor (Csengeri et al., 2017) with a mass of that hosts the MYSO. For this reason, the source is included in the ALMAGAL survey, a large program on ALMA (2019.1.00195.L, PI: S. Molinari), dedicated to studying the evolution of high-mass protocluster formation in the Galaxy. Araya et al. (2005) found blueshifted and redshifted CS (3–2) as well as 13CO (2–1) line wings, indicative of a molecular outflow. A similar signature in the 18CO (2–1) line is observed by the SEDIGISM survey (Yang et al., 2022). However, Guerra-Varas et al. (2023) did not detect an outflow in the SiO (2-1) line and wings of the HCO+ (2-1) line with the APEX telescope. Recently, Ma et al. (2023) identified this clump as a hub for star formation fed by three filaments. For the hub junction, they derived an extent of 2.4 pc 2.4 pc and a mass of .
@@
Associated weak radio continuum emission was first reported by Haynes et al. (1978) and later mapped by Walsh et al. (1998) who did not resolve the source (). Araya et al. (2005) detected broad radio recombination lines toward this region, indicating the presence of a hypercompact Hii region (HCHii). Similar measurements of Murphy et al. (2010), Kim et al. (2018) confirmed this finding. Murphy et al. (2010) also estimated the extinction toward G323, based on the depth of the silicate absorption feature in a low-resolution IRAS spectrum, which amounts to mag. The object was observed in the ATOMS survey (ALMA Three-millimeter Observations of Massive Star-forming regions, Liu et al., 2020b) in Autumn 2019. The radio recombination line and the continuum data at 3 mm are used by Zhang et al. (2023) to characterize the HCHii.
@@
The massive star-forming region G323 is associated with methanol, water, and hydroxyl masers. The Class II methanol 6.7 GHz transition was first observed by MacLeod et al. (1992). Monitoring of this transition at Hartebeesthoek Radio Astronomy Observatory (HartRAO) revealed an increase in the total flux density from 20 Jy in 2011 (Green et al., 2015) to 7 500 Jy in 2015 (Proven-Adzri et al., 2019), that is, by a factor of 430. Interestingly, the decay of the methanol maser, which has been ongoing since then, is characterized by a periodic flux variation with a cycle time of 93 d (MacLeod et al., 2021). The first flare evidence was obtained from the brightening of the 6.035 GHz exOH maser (MacLeod et al., 2021). The discovery of the maser flare raises the question whether it is due to an accretion burst, similar to S255IR-NIRS3 (Caratti o Garatti et al., 2017), NGC6334I-MM1 (Hunter et al., 2017, 2018), and G358 (Brogan et al., 2019; MacLeod et al., 2019; Burns et al., 2020; Stecklum et al., 2021). The source is classified as ”irregular” in the recent NEOWISE-based variability study of 6.7 GHz maser sources by Song et al. (2023). Their exclusion of the pre-burst epochs prevents the burst detection.
@@ 2.1 VVV(X) survey imaging
The VISTA Variables in Via Lactea Survey (VVV, Minniti et al., 2010) and its extension VVVX (from 2016 to 2019) are ESO public surveys that were conducted with the 4-m VISTA telescope in the red optical/near-infrared (NIR) domain. They target the Galactic bulge and part of the adjacent plane. Images of the G323 region were retrieved from the VISTA Science Archive. VVV(X) obtained -band imaging in all of their observing epochs of the region. Data for the and filters are available only for 2010 and one epoch in 2015. Although this does not suffice to produce light curves, a possible color change resulting from the burst can be studied nevertheless. NIR images show the bright counterpart of G323 embedded in a scattering nebulosity. Photometry was established on VVV(X) images, taking into account the core saturation of its point spread function (PSF) in the -band. The affected pixels were given zero weight when fitting the image profile. This was performed using the MPFit2DPeak function of the IDL Astronomy Users Library (Landsman, 1995), using a tilted Moffat function that is appropriate for PSFs in ground-based images (Moffat, 1969). Images with seeing worse than 35 were discarded. The photometry is given in Table 10 of the Appendix.
@@ 2.3 ISAAC spectroscopy
Long-slit spectroscopy with the Infrared Spectrometer and Array Camera ISAAC was performed on June 11, 2013, at the ESO-VLT UT3 telescope, Paranal, Chile. A slit of 03 width was used, producing a spectral resolution of 8 900, at a position angle (PA) of 438, centered on the source (see Fig. 10). The total integration time was 4 min. Nodding along the slit was performed with a throw of 35″ to allow for sky subtraction. The grating was set at a central wavelength of 2.15 m, offering a spectral bandwidth of 0.122 m, covering a wavelength range from 2.086 to 2.210 m. Data reduction was performed using standard IRAF222IRAF (Image Reduction and Analysis Facility) is distributed by the National Optical Astronomy Observatories, which are operated by AURA, Inc., in cooperative agreement with the National Science Foundation. tasks. Each observation was flat-fielded, sky subtracted, and corrected for the distortion caused by long-slit spectroscopy. The atmospheric response was corrected by dividing each spectrum by a telluric standard star (acquired with the same science settings), normalized to the blackbody function at the stellar temperature, and corrected for any absorption feature intrinsic to the star. Originally, we used the telluric standard star for flux calibration. However, probably due to slit flux loss, the resulting calibrated spectrum is = 6.73, mag, that is, 0.6, mag lower than the VVV photometry around that epoch (see Sect. 3.2). Therefore, for calibration, we adopted the closest -band photometric point (16 June 2013). Wavelength calibration was performed using the many OH lines located in this spectral range, resulting in an average uncertainty of 0.2 Å in the extracted spectra. Three spectra, centered on the source and a few arcseconds NE and SW off the source, were extracted from the spectral image, where both continuum and line emission are present. Continuum emission from the final spectral image was removed with the IRAF task ”continuum”. Finally, the radial velocities of the observed lines were calculated using single or multiple Gaussian fits (for multiple components), and corrected for the target velocity with respect to the LSR.
@@ 2.9 ALMA observations
The ALMA Cycle 7 project ALMAGAL aims to observe the 1.3 mm continuum and lines toward dense molecular clumps in the Galactic Plane at a sensitivity level of . Measurements are made with the 12-m array and the 7-m Alma Compact Array (ACA). Data for G323 were retrieved from the ALMA Science Archive (e.g., Stoehr et al., 2014). For this paper, we used the available 12-m array measurements on ALMAGAL field 767784 (observing date: 30-December-2019), where we worked on the data products obtained by the standard pipeline in CASA (Common Astronomy Software Applications package, McMullin et al., 2007). This data set has a typical spatial resolution of 12. In the course of this work, a discrepancy was observed between the flux density at 1.4 mm obtained from the ALMAGAL data and the 3 mm measurement by Zhang et al. (2023). Therefore, the corresponding data set from Zhang et al. (2023), obtained at about the same epoch, was also analyzed using CASA.
@@ 3.3 The scattered light echo
Light echoes (LEs) are also detected in the NEOWISE and images, which were retrieved using the ICORE tool (Masci, 2013). This is not self-evident, as the scattering cross-section of interstellar dust grains at longer wavelengths is smaller (e.g., Draine, 2003). However, the lower scattering efficiency is more than compensated by higher flux densities in the NEOWISE bands compared to (cf. Fig. 15). The features of the LEs and their evolution seen by NEOWISE correspond to what is found in the -band.
Near the source, the spatial brightness distribution of the LE echo represents a record of the burst history. Thus, the maximum of the ratio values provides an estimate of the maximum burst strength. This holds for the - and -bands, where the maximum values at various locations are around 20. Further away, the echo strength decreases as a result of spatial propagation and superposition along the line of sight. This is already the case for the -band. Clearly, the LE is not circular-symmetric but extends southeast and is missing in the opposite direction. This indicates a nonuniform dust distribution in the environment of G323. The absence of an LE in the northwest could be related to the curved rim-shaped structure in this area (cf. Fig. 1) that could block or shadow it.
@@
The remote LE is still faintly visible in NEOWISE images taken in 2023 (cf. Fig. 24). Its light curves (Fig. 7) were established by integrating the fluxes in the , , and bands over this region, marked in Fig. 6. The subtracted background was derived from an area northwest of the MYSO, devoid of stars. Notably, these light curves cover the burst rise which is missing in the on-source data (cf. Fig. 4) because of the time gap between the and NEOWISE missions. The primary characteristics that can be drawn from them are the following. The LE peak occurs slightly more than four years after the burst peak. Its amplitude is shallower than that of the burst, which is due to echo propagation. The color change confirms the ‘bluening’ during the burst. Due to the wavelength dependence of the scattering cross section, the pre-burst colors of the LE are bluer than those of the source, as indicated by vs. 0.12 and vs. .
@@
For common cases of LEs, as for example those associated with novae, supernovae, and variable stars (Sugerman, 2007; Jiang et al., 2016; Kervella et al., 2008), a superluminal motion of the echo is present. It is a pure geometrical effect that follows from first principles in the case of single scattering (Nemiroff et al., 2016). Whether a prompt YSO LE is superluminal is questionable, since its environment is quite dusty. Thus, multiple scattering may overwhelm the geometric effect on which the superluminal motion rests. For G323, the apparent propagation velocity can be derived from the projected separation between the echo boundary and the MYSO, as well as the time difference between the date of the burst peak and the epoch of observation. For the -band, the respective quantities are 22″ and one year. Together with a distance of 4.08 kpc, this yields 1.4 . Although this confirms the superluminal motion in the -band, it implies that the apparent velocity is smaller for and . Given the small extent of the echo, its apparent velocity is probably close to subluminal.
@@
In the optically thick regions, the photons follow random walk paths, where the effective travel distance is given by the diffusion path length. For random walk processes, the diffusion path length is (assuming isotropic scattering), where is the mean free path and is the number of scattering events. Within a fixed time the photons can travel a path with a ”real” length of , with c being the speed of light. The ”real” path is the same for all wavelengths (it depends only on the travel time and c), but the interaction free path length (and hence the number of scattering events) depends on the optical depth (and hence the wavelength). When the equations are combined, the diffusion length at a given time and wavelength is . This implies a difference in the diffusion path length (and hence the apparent LE extent) for different wavelengths. With (see e.g., Krieger & Wolf, 2021), the expected extent of the LE in compared to that seen in Z can be written as the inverse ratio of the optical depths . For MRN dust (Mathis et al., 1977), the optical depth ratio between the Z and filters is . This would imply that LE propagation is almost a factor four slower in Z than in . However, the density within the cloud core is highly variable, and a measurable slowdown is expected only in the densest regions (close to the protostar) and therefore the effect will be smaller.